# TimeFilter: Patch-Specific Spatial-Temporal Graph Filtration for Time Series Forecasting

### v1

@@ Abstract
Current time series forecasting methods can be broadly classified into two categories: Channel Independent (CI) and Channel Dependent (CD) strategies, both aiming to capture the complex dependencies within time series data. However, the CI strategy fails to exploit highly correlated covariate information, while the CD strategy integrates all dependencies, including irrelevant or noisy ones, thus compromising generalization. To mitigate these issues, recent works have introduced the Channel Clustering (CC) strategy by grouping channels with similar characteristics and applying different modeling techniques to each cluster. However, coarse-grained clustering cannot flexibly capture complex, time-varying interactions. Addressing the above challenges, we propose TimeFilter, a graph-based framework for adaptive and fine-grained dependency modeling. Specifically, after constructing the graph with the input sequence, TimeFilter filters out irrelevant correlations and preserves the most critical ones through patch-specific filtering. Extensive experiments on 13 real-world datasets from various application domains demonstrate the state-of-the-art performance of TimeFilter. The code is available at https://github.com/TROUBADOUR000/TimeFilter.
@@ 1 Introduction
Figure 1: Analysis of three channels from the Electricity dataset shows the pros and cons of CI, CD, and CC strategies in different cases. The CI strategy ignores the highly correlated covariate information in the Supporting Case (left). The CD strategy fuses all information, including the irrelevant dependencies in the Supporting Case (right). The CC strategy addresses these issues by using global correlation to form CD and CI clusters. However, it cannot model the dynamic interactions of channels at different time steps at a fine-grained level. For example, in the Failure Case, it still faces the same dilemma as CD and CI strategies.
@@
  • Paradigm Transformation. We for the first time advocate for a fine-grained segmentation of dependencies, capturing spatial-temporal relationships at the patch level while adaptively filtering out irrelevant information.
@@
  • Holistic Evaluation. Comprehensive experiments demonstrate that TimeFilter achieves state-of-the-art performance in both long- and short-term forecasting, with error reductions of $4.48\%$ and $8.02\%$, respectively.
@@ 2.2 Dependency Modeling
Regarding modeling dependencies in MTSF, some models adopt a Channel Independence (CI) strategy (Nie et al., 2023; Zeng et al., 2023; Lin et al., 2024), as shown in Figure 2(a), preserving only the temporal dependencies. These models use only the historical information of each individual sequence, ignoring the interactions between variables. While this approach is robust, it wastes the potentially relevant relationships from inter-channel information. In contrast, the Channel Dependency (CD) strategy (Liu et al., 2024b; Zhang & Yan, 2023; Yu et al., 2024), as shown in Figure 2(b), models the full spectrum of information, providing information gains but also including redundant dependencies, thus reducing robustness. Designing an adaptive Dependency Modeling strategy remains a significant challenge. Recent research (Chen et al., 2024; Qiu et al., 2024b) employs a channel clustering strategy to dynamically select modeling strategies based on channel similarity, as shown in Figure 2(d). For instance, variables $x_{1}$ and $x_{3}$ are chosen to form a CD cluster, where inter-channel relationships are modeled, while variable $x_{2}$ is assigned to a CI cluster, relying only on its own historical values. However, this coarse-grained approach fails to capture the dynamically evolving interactions. Therefore, as shown in Figure 2(c), we design a robust and fine-grained filtering mechanism to customize the dependencies needed for each time segment.
@@ 4.2 Patch-Specific Filtration (PSF) Module
  • Spa.-tem. Filter: $\mathbf{M}_{i}\sim p(\mathbf{M}_{i}|\mathbf{\mathcal{G}}_{i}^{ST})$ If the selection includes a Spatial-temporal Filter, the spatial edges \$\mathbf{\mathcal{E}}\backslash(\mathbf{\mathcal{E}}^{S}\cup\mathbf{\mathcal{E}}_{i}^{T})$ are retained. For instance, in traffic flow prediction, spatial-temporal dependencies are pivotal (Wang et al., 2024), capturing the nuanced interactions between traffic dynamics across different locations and their evolution over time.
Following the classic concept of a (sparsely-gated) mixture-of-experts (Shazeer et al., 2017b; Zhang et al., 2024), we calculate the confidence $r$ of the current patch for each filter using a noisy gating mechanism.
@@ 5.1 Experimental Setup
For long-term forecasting, we conduct extensive experiments on selected $9$ commonly real-world datasets from various application scenarios, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Traffic, Electricity, Weather, Solar-Energy and Climate (with $1763$ channels) datasets. For short-term forecasting, we adopt $4$ benchmarks from PEMS (PEMS03, PEMS04, PEMS07, PEMS08). The statistics of the dataset are shown in Sec. A.1.
@@ 5.2 Results
Models TimeFilter iTransformer Leddam MSGNet SOFTS CrossGNN PatchTST Crossformer TimesNet DLinear FEDformer
@@
ETT (Avg.) 0.358 0.385 0.383 0.399 0.368 0.388 0.384 0.403 0.379 0.396 0.371 0.392 0.381 0.397 0.685 0.578 0.391 0.404 0.446 0.447 0.408 0.428
Weather 0.239 0.269 0.258 0.279 0.242 0.272 0.249 0.278 0.255 0.278 0.247 0.289 0.259 0.281 0.259 0.315 0.259 0.287 0.265 0.315 0.309 0.360
Electricity 0.158 0.256 0.178 0.270 0.169 0.263 0.194 0.300 0.174 0.264 0.201 0.300 0.216 0.304 0.244 0.334 0.193 0.295 0.225 0.319 0.214 0.327
Traffic 0.407 0.268 0.428 0.282 0.467 0.294 0.641 0.370 0.409 0.267 0.583 0.323 0.555 0.362 0.550 0.304 0.620 0.336 0.625 0.383 0.610 0.376
Solar-Energy 0.223 0.250 0.233 0.262 0.230 0.264 0.262 0.288 0.229 0.256 0.324 0.377 0.270 0.307 0.641 0.639 0.301 0.319 0.330 0.401 0.292 0.381
Table 1: Long-term forecasting results. The input length $L$ is $96$. All results are averaged across four different forecasting horizon: $T\in\{96,192,336,720\}$. The best and second-best results are highlighted in bold and underlined, respectively. See Tab. 8 for full results.
@@
Table 2: Long-term forecasting results. The input length $L$ is searched from $\{192,336,512,720\}$ for optimal horizon in the scaling law of TSF (Shi et al., 2024). All results are averaged across four different forecasting horizon: $T\in\{96,192,336,720\}$. The best and second-best results are highlighted in bold and underlined, respectively. See Tab. 9 for full results.
We report the results of the long-term forecasting in Tab. 1. The input length $L$ is $96$ for our method and all baselines. The forecasting horizon $T$ is $\{96,192,336,720\}$. TimeFilter adaptively models various dependencies for each dataset in different domains, and has been proven to yield significant improvements. From the table, compared to the second-best model, Leddam (Yu et al., 2024), TimeFilter reduces MSE and MAE by $4.48\%/2.23\%$, respectively.
Furthermore, according to the scaling law (Shi et al., 2024) of TSF, for a given dataset size, as the look-back horizon $L$ increases, so does the noise in the input sequence, which does not necessarily improve the performance of the model. In addition, the dataset size has more influence on the optimal horizon, while the model size has less influence. Therefore, we conduct a search for different look-back horizons to determine the optimal horizon for the $4$ relatively large datasets and compare TimeFilter with other baselines. As shown in Tab. 2, TimeFilter demonstrates robust resistance to noise and maintains SOTA performance even with longer optimal horizon inputs, outperforming other methods. Specifically, compared to the channel clustering strategy methods DUET (Qiu et al., 2024b) and CCM (Chen et al., 2024), TimeFilter reduces MSE and MAE by $5.34\%/1.40\%$ and $6.89\%/3.69\%$, respectively. This demonstrates the effectiveness of our fine-grained adaptive filtration strategy.
Short-term Forecasting.
We report the results of the short-term forecasting in Tab. 3. The input length $L$ is $96$ for our method and all baselines. The forecasting horizon $T$ is $\{12,24,48\}$. From the table, TimeFilter consistently outperforms other methods across all $4$ PEMS datasets. Specifically, on PEMS08, TimeFilter reduces MSE and MAE by $18.63\%/11.85\%$ compared to the second-best model, Leddam (Yu et al., 2024). On PEMS07 with $883$ channels, TimeFilter reduces MSE and MAE by $5.33\%/1.73\%$ compared to the second-best model, SOFTS (Han et al., 2024a).
Models TimeFilter iTransformer Leddam SOFTS PatchTST Crossformer TimesNet TiDE DLinear SCINet FEDformer
@@
PEMS03 0.084 0.191 0.096 0.204 0.101 0.210 0.087 0.192 0.151 0.265 0.138 0.253 0.119 0.271 0.271 0.380 0.219 0.295 0.093 0.203 0.167 0.291
PEMS04 0.083 0.186 0.098 0.207 0.102 0.213 0.091 0.196 0.162 0.273 0.145 0.267 0.109 0.220 0.307 0.405 0.236 0.350 0.085 0.194 0.195 0.308
PEMS07 0.071 0.170 0.088 0.190 0.087 0.192 0.075 0.173 0.166 0.270 0.181 0.272 0.106 0.208 0.297 0.394 0.241 0.343 0.112 0.211 0.133 0.282
PEMS08 0.083 0.186 0.127 0.212 0.102 0.211 0.114 0.208 0.238 0.289 0.232 0.270 0.150 0.244 0.347 0.421 0.281 0.366 0.133 0.225 0.234 0.326
Table 3: Short-term forecasting results. The input length $L$ is $96$. All results are averaged across three different forecasting horizon: $T\in\{12,24,48\}$. The best and second-best results are highlighted in bold and underlined, respectively. See Tab. 10 for full results.
@@ 5.3 Ablation Study
Table 4: Ablation study of TimeFilter. The look-back horizon $L$ is $96$. See Tab. 11 for full results.
@@
We carefully design six other filtering methods and the comparison is shown in the Filter Replace section of Tab. 4. Top-$K$ strategy selects the top $K$ edges with the largest weights from the ego graph $\mathbf{\mathcal{G}}_{i}$ of each patch, while Random-$K$ randomly selects $K$ edges. RegionTop-$K$ strategy selects the top $K$ edges with the largest weights within the region corresponding to three types of dependencies, while RegionThre learns a threshold for each region through a linear mapping and selects edges with weights greater than the threshold in each region. C-Filter refers to channel-wise filtering, where all patches within the look-back sequence are filtered based on the same type of dependency. The w/o Filter refers to the case where no filtering is applied after graph construction.
@@ 6 Conclusion
Considering the complex dynamic interactions inherent in real-world multivariate time series, we propose TimeFilter to adaptively, finely, and robustly model the dependencies. Technically, we design patch-specific filtration to effectively remove redundant relationships, thereby exploiting the most important dependencies for accurate forecasting. Extensive experiments show that TimeFilter consistently achieves state-of-the-art performance in both long- and short-term forecasting tasks. Our work provides an in-depth exploration of dependency modeling in time series forecasting, which we believe may pave the way for further research into effective dependency representations.
@@ A.1 Datasets
We conduct extensive experiments on eight widely-used time series datasets for long-term forecasting and PEMS datasets for short-term forecasting. We report the statistics in Tab. 5. Detailed descriptions of these datasets are as follows:
@@ A.3 Experiment details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on an NVIDIA A100 40GB GPU. We use the Adam optimizer (Kingma, 2014). The batch size is set to 16 for the Electricity and Traffic datasets, and 32 for all other datasets. Tab. 6 provides detailed hyperparameter settings for each dataset.
@@ B.1 Error Bars
In this paper, we repeat all the experiments three times. Here we report the standard deviation of our proposed TimeFilter and the second best model Leddam (Yu et al., 2024), as well as the statistical significance test in Tab. 7. We perform the Wilcoxon test with Leddam (Yu et al., 2024) and obtain the p-value of $4.66e^{-10}$, indicating a significant improvement at the $99\%$ confidence level.
@@ B.2 Long-term Forecasting
Tab. 8 and Tab. 9 present the full results for long-term forecasting, including both the results for the fixed look-back horizon $L=96$ results and the results obtained through hyperparameter search for the optimal horizon according to the scaling law (Shi et al., 2024) of TSF. The hyperparameter search process involved exploring look-back horizons $L\in\{192,336,512,720\}$. In both settings, TimeFilter consistently achieves the best performance, demonstrating its effectiveness and robustness. In particular, under the optimal horizon condition, TimeFilter can still adaptively model effective dependencies in long sequences.
Models TimeFilter iTransformer Leddam MSGNet SOFTS CrossGNN PatchTST Crossformer TimesNet DLinear FEDformer

### v2

@@ Abstract
Time series forecasting methods generally fall into two main categories: Channel Independent (CI) and Channel Dependent (CD) strategies. While CI overlooks important covariate relationships, CD captures all dependencies without distinction, introducing noise and reducing generalization. Recent advances in Channel Clustering (CC) aim to refine dependency modeling by grouping channels with similar characteristics and applying tailored modeling techniques. However, coarse-grained clustering struggles to capture complex, time-varying interactions effectively. To address these challenges, we propose TimeFilter, a GNN-based framework for adaptive and fine-grained dependency modeling. After constructing the graph from the input sequence, TimeFilter refines the learned spatial-temporal dependencies by filtering out irrelevant correlations while preserving the most critical ones in a patch-specific manner. Extensive experiments on 13 real-world datasets from diverse application domains demonstrate the state-of-the-art performance of TimeFilter. The code is available at https://github.com/TROUBADOUR000/TimeFilter.
Machine Learning, ICML
@@ 1 Introduction
Figure 1: Analysis of three channels from the Electricity dataset shows the pros and cons of CI, CD, and CC strategies in different cases. Dynamic Time Warping (DTW) is a metric for measuring the similarity between two sequences, with lower values indicating higher similarity. The CI strategy ignores the highly correlated covariate information in the Supporting Case (left). The CD strategy fuses all information, including the irrelevant dependencies in the Supporting Case (right). The CC strategy addresses these issues by using global correlation to form CD and CI clusters. However, it cannot model the dynamic interactions of channels at different time steps at a fine-grained level. For example, in the Failure Case, it still faces the same dilemma as CD and CI strategies.
@@
  • Novel Paradigm. We for the first time advocate for a fine-grained segmentation of dependencies, capturing spatial-temporal relationships at the patch level while adaptively filtering out irrelevant information.
@@
  • Holistic Evaluation. Comprehensive experiments demonstrate that TimeFilter achieves state-of-the-art performance in both long- and short-term forecasting, with error reductions of $4.48\%$ and $5.34\%$, respectively.
@@ 2.2 Dependency Modeling
Regarding modeling dependencies in MTSF, some models adopt a Channel Independence (CI) strategy (Nie et al., 2023; Zeng et al., 2023; Lin et al., 2024), as in Figure 2(a), preserving only the temporal dependencies. These models use only the historical information of each individual sequence, ignoring the interactions between variables. Although the CI approach is robust, it wastes the potentially relevant relationships from inter-channel information. In contrast, the Channel Dependency (CD) strategy (Liu et al., 2024b; Zhang & Yan, 2023; Yu et al., 2024), as in Figure 2(b), models the full spectrum of information, providing information gains but also including redundant dependencies, thus reducing robustness. Therefore, designing an adaptive dependency modeling strategy remains a significant challenge. Recent research (Chen et al., 2024; Qiu et al., 2024b) employs a Channel Clustering (CC) strategy to dynamically select modeling strategies based on channel similarity, as in Figure 2(d). For instance, variables $x_{1}$ and $x_{3}$ are chosen to form a CD cluster, where inter-channel relationships are modeled, while variable $x_{2}$ is assigned to a CI cluster, relying only on its own historical values. However, this coarse-grained approach fails to capture the dynamically evolving interactions. Therefore, as in Figure 2(c), we design a robust and fine-grained filtering mechanism to customize the dependencies needed for each time segment.
@@ 4.2 Patch-Specific Filtration (PSF) Module
  • Spa.-tem. Filter: $\mathbf{M}_{i}\sim p(\mathbf{M}_{i}|\mathbf{\mathcal{G}}_{i}^{ST})$ If the selection includes a Spatial-temporal Filter, the spatial edges \$\mathbf{\mathcal{E}}\backslash(\mathbf{\mathcal{E}}^{S}\cup\mathbf{\mathcal{E}}_{i}^{T})$ are retained. For instance, in traffic flow prediction, spatial-temporal dependencies are pivotal (Wang et al., 2024b), capturing the nuanced interactions between traffic dynamics across different locations and their evolution over time.
Following the classic concept of a (sparsely-gated) mixture-of-experts (Shazeer et al., 2017b; Zhang et al., 2024a) and graph sparsification paradigms (Wang et al., 2024a; Zhang et al., 2024b, c), we calculate the confidence $r$ of the current patch for each filter using a noisy gating mechanism.
@@ 5.1 Experimental Setup
For long-term forecasting, we conduct extensive experiments on selected $9$ commonly real-world datasets from various application scenarios, including ETT (ETTh1, ETTh2, ETTm1, ETTm2), Traffic, Electricity, Weather, Solar-Energy (Miao et al., 2024a, b) and Climate (with $1763$ channels) datasets. For short-term forecasting, we adopt $4$ benchmarks from PEMS (PEMS03, PEMS04, PEMS07, PEMS08) (Liu et al., 2025b, a). The statistics of the dataset are shown in Sec. A.1.
@@ 5.2 Results
Models TimeFilter DUET iTransformer Leddam MSGNet SOFTS CrossGNN PatchTST Crossformer TimesNet DLinear FEDformer
CC (2024b)
@@
MSE
MAE
ETT (Avg.) 0.358 0.385 0.371 0.388 0.383 0.399 0.368 0.388 0.384 0.403 0.379 0.396 0.371 0.392 0.381 0.397 0.685 0.578 0.391 0.404 0.446 0.447 0.408 0.428
Weather 0.239 0.269 0.251 0.273 0.258 0.279 0.242 0.272 0.249 0.278 0.255 0.278 0.247 0.289 0.259 0.281 0.259 0.315 0.259 0.287 0.265 0.315 0.309 0.360
Electricity 0.158 0.256 0.172 0.258 0.178 0.270 0.169 0.263 0.194 0.300 0.174 0.264 0.201 0.300 0.216 0.304 0.244 0.334 0.193 0.295 0.225 0.319 0.214 0.327
Traffic 0.407 0.268 0.451 0.269 0.428 0.282 0.467 0.294 0.641 0.370 0.409 0.267 0.583 0.323 0.555 0.362 0.550 0.304 0.620 0.336 0.625 0.383 0.610 0.376
Solar-Energy 0.223 0.250 0.237 0.233 0.233 0.262 0.230 0.264 0.262 0.288 0.229 0.256 0.324 0.377 0.270 0.307 0.641 0.639 0.301 0.319 0.330 0.401 0.292 0.381
Table 1: Long-term forecasting results. The input length $L$ is $96$. All results are averaged across four different forecasting horizons: $T\in\{96,192,336,720\}$. The best and second-best results are in bold and underlined, respectively. See Table 8 for full results.
@@
Table 2: Long-term forecasting results. The input length $L$ is searched from $\{192,336,512,720\}$ for optimal horizon in the scaling law of TSF (Shi et al., 2024). All results are averaged across four different forecasting horizons: $T\in\{96,192,336,720\}$. The best and second-best results are in bold and underlined, respectively. See Table 9 for full results.
Models TimeFilter DUET iTransformer Leddam SOFTS PatchTST Crossformer TimesNet TiDE DLinear SCINet FEDformer
CC (2024b)
@@
MSE
MAE
PEMS03 0.084 0.191 0.086 0.192 0.096 0.204 0.101 0.210 0.087 0.192 0.151 0.265 0.138 0.253 0.119 0.271 0.271 0.380 0.219 0.295 0.093 0.203 0.167 0.291
PEMS04 0.083 0.186 0.096 0.203 0.098 0.207 0.102 0.213 0.091 0.196 0.162 0.273 0.145 0.267 0.109 0.220 0.307 0.405 0.236 0.350 0.085 0.194 0.195 0.308
PEMS07 0.071 0.170 0.076 0.176 0.088 0.190 0.087 0.192 0.075 0.173 0.166 0.270 0.181 0.272 0.106 0.208 0.297 0.394 0.241 0.343 0.112 0.211 0.133 0.282
PEMS08 0.083 0.186 0.096 0.192 0.127 0.212 0.102 0.211 0.114 0.208 0.238 0.289 0.232 0.270 0.150 0.244 0.347 0.421 0.281 0.366 0.133 0.225 0.234 0.326
Table 3: Short-term forecasting results. The input length $L$ is $96$. All results are averaged across three different forecasting horizons: $T\in\{12,24,48\}$. The best and second-best results are in bold and underlined, respectively. See Table 10 for full results.
We report the results of the long-term forecasting in Table 1. The input length $L$ is $96$ for our method and all baselines. The forecasting horizon $T$ is $\{96,192,336,720\}$. The results show that TimeFilter adaptively models various dependencies for each dataset in different domains, and has been proven to yield significant improvements. Compared to the second-best model Leddam (Yu et al., 2024), TimeFilter reduces MSE and MAE by $4.48\%/2.23\%$, respectively.
Furthermore, according to the scaling law (Shi et al., 2024) of TSF, for a given dataset size, as the look-back horizon $L$ increases, so does the noise in the input sequence, which does not necessarily improve the performance of the model. In addition, the dataset size has more influence on the optimal horizon, while the model size has less influence. Therefore, we conduct a search for different look-back horizons to determine the optimal horizon for the $4$ relatively large datasets and compare TimeFilter with other baselines. As shown in Table 2, TimeFilter demonstrates robust resistance to noise and maintains SOTA performance even with longer optimal horizon inputs, outperforming other methods. Specifically, compared to the channel clustering strategy methods DUET (Qiu et al., 2024b) and CCM (Chen et al., 2024), TimeFilter reduces MSE and MAE by $5.34\%/1.40\%$ and $6.89\%/3.69\%$, respectively. This demonstrates the effectiveness of our fine-grained adaptive filtration strategy.
Short-term Forecasting.
We report the results of the short-term forecasting in Table 3. The input length $L$ is $96$ for our method and all baselines. The forecasting horizon $T$ is $\{12,24,48\}$. From the table, TimeFilter consistently outperforms other methods across all $4$ PEMS datasets. Specifically, on PEMS08, TimeFilter reduces MSE and MAE by $13.54\%/3.13\%$ compared to the second-best DUET (Qiu et al., 2024b). On PEMS07 with $883$ channels, TimeFilter reduces MSE and MAE by $5.33\%/1.73\%$ compared to the second-best SOFTS (Han et al., 2024a).
@@ 5.3 Ablation Study
Table 4: Ablation study of TimeFilter. The The input length $L$ is $96$. See Table 11 for full results.
@@
We carefully design six other filtering methods and the comparison is shown in the Filter Replace section of Table 4. Top-$K$ strategy selects the top $K$ edges with the largest weights from the ego graph $\mathbf{\mathcal{G}}_{i}$ of each patch, while Random-$K$ randomly selects $K$ edges. RegionTop-$K$ strategy selects the top $K$ edges with the largest weights within the region corresponding to three types of dependencies, while RegionThre learns a threshold for each region through a linear mapping and selects edges with weights greater than the threshold in each region. C-Filter refers to channel-wise filtering, where all patches within the look-back sequence are filtered based on the same type of dependency. The w/o Filter refers to the case where no filtering is applied after graph construction.
@@ 6 Conclusion
Considering the complex dynamic interactions inherent in real-world multivariate time series, we propose TimeFilter to adaptively, finely, and robustly model the dependencies. Technically, we design patch-specific filtration to effectively remove redundant relationships, thereby exploiting the most important dependencies for accurate forecasting. Extensive experiments show that TimeFilter consistently achieves SOTA performance in both long- and short-term forecasting tasks. Our work provides an in-depth exploration of dependency modeling in TSF, which we believe may pave the way for further research into dependency representations.
Acknowledgements
Dai Tao is supported in part by the National Natural Science Foundation of China, under Grant (62302309, 62171248), Shenzhen Science and Technology Program (JCYJ20220818101014030, JCYJ20220818101012025).
Impact Statement
Time series forecasting plays a crucial role in a variety of real-world domains, such as finance, weather forecasting, and traffic control. Our study presents a novel method for enhancing the dependency modeling. The datasets utilized are publicly accessible, promoting transparency and reproducibility. We do not anticipate any ethical issues arising from this work. The primary contributions of this research advance the field of time series forecasting, with potential applications in multiple sectors.
@@ A.1 Datasets
We conduct extensive experiments on eight widely-used time series datasets for long-term forecasting and PEMS datasets for short-term forecasting. We report the statistics in Table 5. Detailed descriptions of these datasets are as follows:
@@ A.3 Experiment details
All experiments are implemented in PyTorch (Paszke et al., 2019) and conducted on an NVIDIA A100 40GB GPU. We use the Adam optimizer (Kingma, 2014). The batch size is set to 16 for the Electricity and Traffic datasets, and 32 for all other datasets. Table 6 provides detailed hyperparameter settings for each dataset.
@@ B.1 Error Bars
In this paper, we repeat all the experiments three times. Here we report the standard deviation of our proposed TimeFilter and the second best model Leddam (Yu et al., 2024), as well as the statistical significance test in Table 7. We perform the Wilcoxon test with Leddam (Yu et al., 2024) and obtain the p-value of $4.66e^{-10}$, indicating a significant improvement at the $99\%$ confidence level.
@@ B.2 Long-term Forecasting
Table 8 and Table 9 present the full results for long-term forecasting, including both the results for the fixed look-back horizon $L=96$ results and the results obtained through hyperparameter search for the optimal horizon according to the scaling law (Shi et al., 2024) of TSF. The hyperparameter search process involved exploring look-back horizons $L\in\{192,336,512,720\}$. In both settings, TimeFilter consistently achieves the best performance, demonstrating its effectiveness and robustness. In particular, under the optimal horizon condition, TimeFilter can still adaptively model effective dependencies in long sequences.
Models TimeFilter DUET iTransformer Leddam MSGNet SOFTS CrossGNN PatchTST Crossformer TimesNet DLinear FEDformer